USGS WaterAlert System
Little Bear Fire Water Alert System USGS Water Alert http://water.usgs.gov/wateralert Subscribe to gages to receive email or text message when water is detected. Please see attached handout for a... more
Little Bear BAER Update - July 31, 2012
Aerial application of grass seed
Aerial application of grass seed was completed as planned on 19,211 acres on Sunday, July 29, 2012.
Aerial application of straw mulch
As of July 30th, aerial application of straw mulch has been completed on 8,748 acres of the 10,311 total acres planned for treatment. The remaining 1,563 acres should be completed by the end of this week. There will be some re-mulching of specific light areas, but straw mulch operations are about 85% complete.
• Shawn Wilson Excavating is still working on FR 127A. All bridges now have a by-pass swale constructed to permit flows greater than the bridges can handle to bypass the bridges. Gabion baskets are being installed at a major cross-drainage, as well as along a stream bed to protect a cabin situated at the edge of the stream.
• FR 600 (Kraut Canyon) is closed for now and no additional work is being done at this time.
• FR 107 (Bonito) was graded and approx. 50 cubic yards of rock was placed in a soft spot in the road at the junction with FR108. The adjacent ditch was also cleaned.
• Rip-rap is being hauled into the area of FR107(C) and South Fork Campground in preparation for fortification in several areas.
• FR 958 (Philadelphia Canyon) has been heavily re-worked to repair damage and prepare for the next storm.
• Work along FR 117 (Monjeau) has been on hold due to traffic from the straw mulch operations, but work along this road will resume soon.
Hazard tree removal is mostly complete, though we may identify and take down a few more trees as we go.
Work to repair and re-establish the trails affected by the Little Bear fire is on hold due to the runoff from the rains. Trails work will be underway once the monsoon rains are over.
US Geological Survey Stream and Precipitation Gages / WaterAlert System
The precipitation and stream gages are working well and feedback is positive.
Lincoln County Watershed Protection Task Force Update (multi-agency flood response effort)
• Removal of woody debris from waterways and drainages is complete.
• Water from Bonito Lake continues to be released downstream using 5 large pumps. The lake level was 28'8" below the spillway as of 7/27/2012. The bacteria level at the lake is very high and the public is encouraged not to come in contact with the lake water. Water samples are also being taken from Rio Ruidoso and Eagle Creek.
• Work continues at Alto Lake/Reservoir to fortify the dam, but the work at the duck pond above Alto Lake is mostly complete. Approximately 70,000 yards of material has been hauled out of the duck pond.
• Upper Hondo Soil and Water Conservation District continues to assemble and provide mulch wattles to the public at no charge. Seed and wood mulch are also available at no charge through their Capitan office. Sand bags are available at no charge at the Lincoln County Office of Emergency Services.
• This task force will demobilize and turn emergency response back over to Lincoln County Office of Emergency Services as of August 3, 2012. However, the task force can be reassembled if needed to respond to a flood-related emergency in the area of Little Bear fire.
|Incident Type||Burned Area Emergency Rehabilitation|
|Date of Origin||Wednesday June 20th, 2012 approx. 12:00 AM|
|Location||Lincoln County, NM|
|Incident Commander||Dave Warnack|
|Estimated Containment Date||Thursday July 05th, 2012 approx. 06:00 PM|